OpenStudy (samigupta8):

Can anyone plz...help me in understanding the concept of non negative integral solution of X1+X2+X3+.....Xn=m ???

Parth (parthkohli):

Oh, this is great. I know a pretty good way to understand this. Hold on.

OpenStudy (anonymous):

pleaseb type full question as you see in your book

Parth (parthkohli):

Imagine five stars like this:\[\large \star \star \star \star \star\]If we want to find the number of solutions to \(x_1 + x_2 + x_3= 5\), we can first think about how to represent solutions graphically. The following is a graphical interpretation of a solution:\[\large \star \star \star |\star| \star \]This represents \(3 + 1 + 1 = 5\). You can represent all solutions this way.

OpenStudy (samigupta8):

This will turn out to be very lengthy method.....doing like dis vil help in getting the ans aftr a gr8 consumption of time...

Parth (parthkohli):

So like\[\large | \star \star| \star \star \star\]The above represents \(0 + 2 + 3 = 5\). And no, we're not going to solve it like this. We're going to derive the expression\[\binom{m+n-1}{n-1}\]

OpenStudy (samigupta8):

This is what i want to ask how do v get this formula ...??

Parth (parthkohli):

That's what I'm doing. So do you get what I'm doing above? Like how I'm graphically representing the solutions If you understand that we'll proceed with the derivation

OpenStudy (samigupta8):

Yup....i got it...

Parth (parthkohli):

OK, so finding the number of non-negative solutions to \(x_1 +x_2+x_3 + \cdots + x_n = m\) is the same as the number of ways to separate \(m\) stars by \(n-1\) bars. Now the interesting part... Think of \(\large \star \star \star |\star| \star\) as a seven letter word. We have to choose two places for the bars and the rest of the places will be taken by the stars. So the number of solutions of \(x_1 + x_2 + x_3 = 5\) is \(\binom{7}2\).

Parth (parthkohli):

Similarly, in general, we have \(m\) stars and \(n-1\) bars. Now we have an \(m+n-1\) letter word out of which we will choose \(n-1\) places for the bars. Therefore the number of solutions is\[\binom{m+n-1}{n-1}\]

Parth (parthkohli):

Do you understand this approach?

OpenStudy (samigupta8):

I m trying to understand it....:)

OpenStudy (samigupta8):

U said that we need to place bars in between the letter words...ryt ...

Parth (parthkohli):

Yeah, you can think of it that way.

OpenStudy (samigupta8):

It is like dat way evn if we were to arrange the stars between the letter wrdd then also we wil get same result bcz then bars will automatically occupy the left over places...

Parth (parthkohli):

\[x_1 + x_2 + x_3 +x_4= 7\]Seven stars, three bars. Total ten "letters". We have to choose three places for the bars. Total 10C3 solutions.